# How To Embed Matters: Evaluation of EO Embedding Design Choices

### 11 Mar 2026 v1

@@ Abstract
Earth observation (EO) missions produce petabytes of multispectral imagery, increasingly analyzed using large Geospatial Foundation Models (GeoFMs). Alongside end-to-end adaptation, workflows make growing use of intermediate representations as task-agnostic embeddings, enabling models to compute representations once and reuse them across downstream tasks. Consequently, when GeoFMs act as feature extractors, decisions about how representations are obtained, aggregated, and combined affect downstream performance and pipeline scalability. Understanding these trade-offs is essential for scalable embedding-based EO workflows, where compact embeddings can replace raw data while remaining broadly useful. We present a systematic analysis of embedding design in GeoFM‑based EO workflows. Leveraging NeuCo‑Bench, we study how backbone architecture, pretraining strategy, representation depth, spatial aggregation, and representation combination influence EO task performance. We demonstrate the usability of GeoFM embeddings by aggregating them into fixed‑size representations more than 500× smaller than the raw input data. Across models, we find consistent trends: transformer backbones with mean pooling provide strong default embeddings, intermediate ResNet layers can outperform final layers, self‑supervised objectives exhibit task‑specific strengths, and combining embeddings from different objectives often improves robustness.
@@ 1 Introduction
In parallel, machine learning for EO has evolved from pixel-level classifiers and hand-crafted features to convolutional and transformer-based architectures that learn high-dimensional feature representations [37, 31]. This shift has led to Geospatial Foundation Models (GeoFMs), pretrained on large-scale EO corpora and adapted to downstream tasks via end-to-end fine-tuning [3, 34]. While effective, such pipelines require repeated access to raw imagery and backbone models, incurring computational and storage costs. To mitigate these limitations, GeoFMs can instead be used to extract generic embeddings that are reused across multiple downstream tasks, offering improved computational efficiency, lightweight decoders, decentralized deployment, and similarity search at scale. In this embedding centric paradigm, the backbone serves as a fixed feature extractor whose representations are computed once and reused across tasks [20]. This reduces the need to repeatedly access and process raw imagery, enabling compact embedding stores that can replace raw data with effective compression ratios, depending on the downstream task and representation design. At EO scale, this supports efficient storage, I/O, and retrieval workflows, particularly when dense spatial or temporal representations are aggregated into compact fixed‑size embeddings tailored to task requirements.
With the rapid release and adoption of new GeoFMs, embeddings derived from these models are becoming central artifacts in EO pipelines. In this setting, design decisions about where representations are extracted, how they are spatially or temporally aggregated, and which information is retained directly shape downstream performance, storage footprint, and scalability. Yet it remains unclear which representation characteristics matter most for different tasks, which aggregation strategies are most effective, and how GeoFM outputs can be transformed into compact, fixed‑size embeddings that remain broadly useful across applications. As a result, principled design of compact and effective EO embeddings remains an open research challenge.
@@ 2 Related Work
EO research increasingly publishes and leverages precomputed vector representations (embeddings) that summarize scene content and enable efficient downstream use [20]. These task‑agnostic embeddings capture general spatial, spectral, or spatiotemporal structure and support broad reuse. Examples include SatCLIP [19], which derives global location embeddings via contrastive image–coordinate alignment; TESSERA [10], which produces pixel-wise temporal embeddings from multi-temporal observations; and AlphaEarth Foundations [4], which models continuous spatiotemporal embedding fields to summarize multi-sensor EO. These efforts advance an embedding centric paradigm in which GeoFMs are trained to produce reusable representations as standalone data products. In parallel, many existing GeoFMs developed for end‑to‑end adaptation are increasingly used as off the shelf feature extractors [6], with the backbone acting as a frozen encoder. However, design choices around representation depth, aggregation, and fixed‑size construction remain largely heuristic. In contrast, we systematically analyze embedding extraction from frozen GeoFMs, examining how architecture, layer selection, pooling, and pretraining objective influence downstream performance and robustness, providing guidance for embedding-centric GeoFM design and use.
@@
Benchmarking GeoFMs typically relies on diverse downstream tasks. Two established GeoFM benchmarks, PANGEA [22] and GEO-BENCH [21, 25], span diverse geospatial domains and task types, including segmentation, classification, change detection, and regression. While differing in design choices, both assume access to the backbone during downstream adaptation allowing downstream models to reuse intermediate feature maps via multi-scale encoder features (e.g., in U-Net-style decoders), rather than constraining the representation to a single fixed-size embedding as in our setting. In contrast, NeuCo-Bench [27], which we use in this work, is a model-agnostic framework for evaluating compact fixed-size embeddings. It comprises per-image regression and classification tasks spanning land cover, biomass, clouds, and heat-island prediction, and is explicitly designed to assess EO embeddings derived from spatiotemporal inputs. Section 3 details how we use NeuCo‑Bench for our embedding design analysis.
@@ 3 Methodology
We study embedding‑centric EO workflows in which pretrained GeoFMs serve as frozen feature extractors, evaluating their representations across diverse downstream tasks. Our analysis focuses on how embedding design choices—encoder family, intermediate layer selection, spatial and temporal aggregation, and representation combination—shape downstream regression performance and robustness. All embeddings are evaluated using the NeuCo‑Bench framework [27].
@@ 3.1 Evaluation Protocol
For each embedding method $p$ and task $t$, we perform $K=50$ repeated random train–test splits. On each split $k$, a linear regressor is trained on the training subset and evaluated on the held-out data, yielding a test-set $R^{2}$ score $s_{k}$. Performance is summarized by the mean across splits,
@@
The public NeuCo-Bench tasks are based on the SSL4EO-S12-downstream dataset [8], which follows the data structure of SSL4EO-S12 [28, 1]. Each input sample consists of four seasonal timesteps with 27 available bands from Sentinel-1 and Sentinel-2 (L2A and L1C), forming a $264\times 264$ patch at 10,m resolution. Each spatiotemporal input cube is paired with a scalar target. We evaluate embeddings on eight regression tasks: Biomass Mean + Std (GEDI‑derived above‑ground biomass statistics), Crops (corn and soybean coverage), Clouds (mean cloud cover), Landcover agricultural + forest (area fractions), and Heat Island Mean + Std (urban surface temperature statistics). Although all tasks are formulated as aggregated regression targets, their underlying signal characteristics differ. We distinguish between semantic proportion tasks, continuous physical measurement tasks, and atmospheric state indicators.
Landcover (Agriculture, Forest) and Crops represent aggregated semantic proportions derived from categorical land-cover maps, where the target reflects the fraction of a patch assigned to discrete classes. Biomass and Heat Island originate from continuous physical measurements—LiDAR-based biomass estimates and thermal surface temperature—that are spatially aggregated to patch-level mean and standard deviation statistics. Clouds, while also modeled as a continuous regression target, is derived from per-acquisition cloud masks and reflects temporally varying atmospheric state rather than a direct surface biophysical property. These differing semantic, structural, and temporal characteristics enable analysis of task-consistent embedding trends across different signal types.
@@ 3.2 GeoFM Backbones and Pretraining Strategies
We evaluate SSL4EO-pretrained ResNet-50 [15] and ViT-Small [7] backbones [28] to maintain comparability in pretraining data. We include TerraMind (ViT-Small) [17] as a reference, as its TerraMind-Base [18] variant achieved leading performance in the evaluations in the original NeuCo-Bench paper. SSL4EO-pretrained backbones are accessed via TorchGeo [26], while TerraMind weights are accessed via TerraTorch [11]. Embeddings are extracted with TerraTorch for consistent processing across models. For the SSL4EO-pretrained backbones, we consider the self-supervised methods DINO [5], MoCo [14], DECUR [29], SoftCon [30], MAE [13], and FGMAE [32]. Not all backbone–method combinations are evaluated due to limited availability in pretrained weights. To ensure comparability, all models are evaluated using Sentinel‑2 L1C inputs only, although TerraMind supports additional modalities.
@@ 3.3 Temporal and Spatial Aggregation
The considered GeoFM encoders produce dense spatial representations: ResNet outputs feature maps of shape $C\times H\times W$, while ViT models produce patch tokens of shape $N_{\text{patch}}\times D$. Image inputs are resized to $224\times 224$ to match pretraining resolution. To obtain a 1D fixed‑size embedding per sample, we apply global mean pooling—across spatial dimensions for ResNet and across tokens for the ViT. We additionally evaluate max and min pooling for both architectures, and for ViTs also analyze CLS‑token embeddings. Embeddings are evaluated at their native dimensionality (2048 for ResNet‑50 and 384 for ViT‑Small).
@@ 4 Experiments and Results
We evaluate embedding design choices for GeoFMs across backbone architecture, pretraining strategy, spatial aggregation, combination and representation depth. We begin with global per-task trends before analyzing each design dimension.
@@ 4.1 General Per-Task Trends
Semantic land-cover targets. Landcover Agriculture and Landcover Forest achieve the highest median $R^{2}$ values (both above 0.80) with low variability, indicating that these signals are reliably captured across embeddings. Crops shows moderate spread. While peak $R^{2}$ values are partially saturated, the Q-score reveals meaningful differences in robustness not visible from mean accuracy alone.
Continuous biophysical and atmospheric targets These targets are more challenging. Biomass and Heat Island Mean show lower median $R^{2}$, and Clouds exhibits strong sensitivity to representation design with wide variation across methods, indicating greater dependence on representation choice. Heat Island Std is the most difficult target overall.
@@
Semantic land‑cover targets. For Crops, Landcover Agriculture, and Landcover Forest, ResNets remain competitive and in some cases slightly exceed ViT performance. This suggests that global semantic composition signals are less sensitive to backbone choice, whereas geophysical and atmospheric variables benefit from richer, long‑range spatial modeling. Across all configurations, TerraMind (ViT‑Small with multimodal pretraining on TerraMesh [2]) achieves the strongest and most stable performance.
These trends align with prior NeuCo‑Bench backbone comparisons performed under matched embedding dimensionality [27], where CNN embeddings were downsampled and ViT embeddings padded. Together with our results, this indicates that the observed ResNet–ViT gap cannot be attributed to embedding dimensionality or resizing artifacts.
@@ 4.3 Task-Specific Effects of SSL Objectives
These trends are consistent across both $R^{2}$ and Q-score. While $R^{2}$ reflects accuracy, the Q-score magnifies robustness differences, and in our results leads to clearer separation between methods. Objective-dependent differences are also more pronounced for transformer backbones. For ResNet models, variation across SSL objectives is weaker, as all methods struggle on continuous biophysical targets and atmospheric targets.
Overall, pretraining objective shapes task-specific representation profiles. Contrastive objectives (e.g., DINO) favor semantic composition signals, whereas reconstruction-based objectives (e.g., MAE/FGMAE) better capture continuous geophysical variation. Rather than yielding a single universally best method, SSL objectives produce complementary strengths, an observation we will revisited in the concatenation experiments.
@@ 4.4 Impact of Spatial Pooling
Figure 3: Per-task Q-Score comparison of spatial aggregation method for ResNet-50 (left) and ViT-Small (right). We use final-layer embeddings with mean, min, or max pooling (or the CLS token for ViT) and average scores across models; negative scores are clipped to zero. For ResNet, mean pooling performs best across tasks, with max pooling outperforming min pooling. For ViT, mean pooling again performs best, with CLS comparable on most tasks, while min and max pooling are similar but weaker—especially on continuous biophysical targets. A corresponding task $R^{2}$ plot is provided in the supplementary material.
@@ 4.5 Complementarity Through Concatenation
We analyze embedding concatenation from two perspectives: (i) per-task gains relative to the stronger task-specific baseline, and (ii) overall gains relative to the strongest baseline across tasks. These two measurements capture distinct effects. When concatenating embeddings under the multi‑task NeuCo‑Bench setting, two effects are possible. First, if individual baselines excel on different tasks, concatenation can improve the overall score simply by retaining the stronger signal per task—even when per‑task gains remain small or negative. Second, if the representations encode genuinely complementary information, concatenation may also yield per‑task improvements beyond both baselines, resulting in gains at both levels.
@@
Across DINO, MAE, and SoftCon, intra‑method concatenation yields only modest improvements (below $\sim$0.03 R2 and below 1 Q‑score point overall), with consistently small per‑task gains (Figure 4). This matches our earlier observation that Mean and CLS embeddings from the same ViT model show highly aligned task profiles and similar overall performance. Given this overlap, little complementary task scores can be exploited for overall improvements.
@@ 4.6 Strength of Intermediate CNN Layers
For ViT‑Small, performance increases over the first few layers and then saturates. Most SSL variants reach near‑peak accuracy by layers 3–5, with only minor changes afterward. Semantic land‑cover targets benefit slightly from deeper layers, whereas continuous biophysical targets and atmospheric targets (Biomass, Clouds, Heat Island) remain relatively stable across depth and, in some cases (e.g., DINO), degrade modestly in the final layers. This suggests that early transformer layers already capture the information needed for aggregated environmental prediction, while deeper layers increasingly emphasize semantic abstraction.
ResNet‑50, in contrast, exhibits a pronounced inverted‑U pattern. Both $R^{2}$ and Q‑score peak at intermediate stages (layers 2–4) and decline at the final stage. The drop is primarily driven by the continuous biophysical targets, which improve substantially at intermediate layers—reaching performance competitive with ViT representations—before degrading in the deepest layer. Semantic land‑cover targets continue to improve more steadily with depth. A resized final‑layer reference confirms that this effect is not attributable to embedding dimensionality. Overall, optimal layer depth is task‑dependent. While final‑layer embeddings are commonly used by default, intermediate layers, especially for CNN backbones, often provide stronger and more stable representations for aggregated biophysical targets. These trends hold consistently across both $R^{2}$ and Q‑score.

### 08 Apr 2026 v2

@@ Abstract
Earth observation (EO) missions produce petabytes of multispectral imagery, increasingly analyzed using large Geospatial Foundation Models (GeoFMs). Alongside end-to-end adaptation, workflows are making growing use of intermediate representations as task-agnostic embeddings, enabling models to compute representations once and reuse them across downstream tasks. Consequently, when GeoFMs act as feature extractors, decisions about how representations are obtained, aggregated, and combined affect downstream performance and pipeline scalability. Understanding these trade-offs is essential for scalable embedding-based EO workflows, where compact embeddings can replace raw data while remaining broadly useful. We present a systematic analysis of embedding design in GeoFM‑based EO workflows. Leveraging NeuCo‑Bench, we study how backbone architecture, pretraining strategy, representation depth, spatial aggregation, and representation combinations influence EO task performance. We demonstrate the usability of GeoFM embeddings by aggregating them into fixed‑size representations more than 500× smaller than the raw input data. Across models, we find consistent trends: transformer backbones with mean pooling provide strong default embeddings, intermediate ResNet layers can outperform final layers, self‑supervised objectives exhibit task‑specific strengths, and combining embeddings from different objectives often improves robustness.
@@ 1 Introduction
In parallel, machine learning for EO has evolved from pixel-level classifiers and handcrafted features to convolutional and transformer-based architectures that learn high-dimensional feature representations [37, 31]. This shift has led to Geospatial Foundation Models (GeoFMs), pretrained on large-scale EO corpora and adapted to downstream tasks through end-to-end fine-tuning [3, 34]. While effective, such pipelines require repeated access to raw imagery and backbone models, incurring computational and storage costs. To mitigate these limitations, GeoFMs can instead be used to extract generic embeddings that are reused across multiple downstream tasks, offering improved computational efficiency, lightweight decoders, decentralized deployment, and similarity search at scale. In this embedding-centric paradigm, the backbone serves as a fixed feature extractor whose representations are computed once and reused across tasks [20]. This reduces the need to repeatedly access and process raw imagery, enabling compact embedding stores that can replace raw data with effective compression ratios, depending on the downstream task and representation design. At EO scale, this supports efficient storage, I/O, and retrieval workflows, particularly when dense spatial or temporal representations are aggregated into compact fixed‑size embeddings.
With the rapid release and adoption of new GeoFMs, embeddings derived from these models are becoming central artifacts in EO pipelines. In this setting, design decisions about where representations are extracted, how they are spatially or temporally aggregated, and which information is retained directly shape downstream performance, storage footprint, and scalability. Yet it remains unclear which representation characteristics matter most for different tasks, which aggregation strategies are most effective, and how GeoFM outputs can be transformed into compact, fixed‑size embeddings that remain broadly useful across applications. As a result, principled design of compact and effective EO embeddings remains an open challenge.
@@ 2 Related Work
EO research increasingly publishes and leverages precomputed vector representations (embeddings) that summarize scene content and enable efficient downstream use [20]. These task‑agnostic embeddings capture general spatial, spectral, or spatiotemporal structure and support broad reuse. Examples include SatCLIP [19], which derives global location embeddings via contrastive image-coordinate alignment; TESSERA [10], which produces pixel-wise temporal embeddings from multi-temporal observations; and AlphaEarth Foundations [4], which models continuous spatiotemporal embedding fields to summarize multi-sensor EO. These efforts advance an embedding-centric paradigm in which GeoFMs are trained to produce reusable representations as standalone data products. In parallel, many existing GeoFMs developed for end‑to‑end adaptation are increasingly used as off-the-shelf feature extractors [6], with the backbone acting as a frozen encoder. However, design choices around representation depth, aggregation, and fixed‑size construction remain largely heuristic. In contrast, we systematically analyze embedding extraction from frozen GeoFMs, examining how architecture, layer selection, pooling, and pretraining objective influence downstream performance and robustness, providing guidance for embedding-centric GeoFM design and use.
@@
Benchmarking GeoFMs typically relies on diverse downstream tasks. Two established GeoFM benchmarks, PANGEA [22] and GEO-BENCH [21, 25], span diverse geospatial domains and task types, including segmentation, classification, change detection, and regression. While differing in design choices, both assume access to the backbone during downstream adaptation allowing downstream models to reuse intermediate feature maps through multi-scale encoder features (e.g., in U-Net-style decoders), rather than constraining the representation to a single fixed-size embedding as in our setting. In contrast, NeuCo-Bench [27], which we use in this work, is a model-agnostic framework for evaluating compact fixed-size embeddings. It comprises per-image regression and classification tasks spanning land cover, biomass, clouds, and heat-island prediction, and is explicitly designed to assess EO embeddings derived from spatiotemporal inputs. Section 3 details how we use NeuCo‑Bench for our embedding design analysis.
@@ 3 Methodology
We study embedding‑centric EO workflows in which pretrained GeoFMs serve as frozen feature extractors, evaluating their representations across diverse downstream tasks. Our analysis focuses on how embedding design choices, including encoder family, intermediate layer selection, spatial and temporal aggregation, and representation combinations, shape downstream regression performance and robustness. All embeddings are evaluated using the NeuCo‑Bench framework [27].
@@ 3.1 Evaluation Protocol
For each embedding method $p$ and task $t$, we perform $K=50$ repeated random train-test splits. On each split $k$, a linear regressor is trained on the training subset and evaluated on the held-out data, yielding a test-set $R^{2}$ score, $s_{k}$. Performance is summarized by the mean across splits,
@@
The public NeuCo-Bench tasks are based on the SSL4EO-S12-downstream dataset [8], which follows the data structure of SSL4EO-S12 [28, 1]. Each input sample consists of four seasonal timesteps with 27 available bands from Sentinel-1 and Sentinel-2 (L2A and L1C), forming a $264\times 264$ patch at 10m resolution. Each spatiotemporal input cube is paired with a scalar target. We evaluate embeddings on eight regression tasks: Biomass Mean + Std (GEDI‑derived above‑ground biomass statistics), Crops (corn and soybean coverage), Clouds (mean cloud cover), Land-cover agricultural + forest (area fractions), and Heat Island Mean + Std (urban surface temperature statistics). Although all tasks are formulated as aggregated regression targets, their underlying signal characteristics differ. We distinguish between semantic proportion tasks, continuous physical measurement tasks, and atmospheric state indicators.
Land-cover (Agriculture, Forest) and Crops represent aggregated semantic proportions derived from categorical land-cover maps, where the target reflects the fraction of a patch assigned to discrete classes. Biomass and Heat Island originate from continuous physical measurements (LiDAR-based biomass estimates and thermal surface temperature) that are spatially aggregated to patch-level mean and standard deviation statistics. Clouds, while also modeled as a continuous regression target, is derived from per-acquisition cloud masks and reflects temporally varying atmospheric state rather than a direct surface biophysical property. These differing semantic, structural, and temporal characteristics enable analysis of task-consistent embedding trends across different signal types.
@@ 3.2 GeoFM Backbones and Pretraining Strategies
We evaluate SSL4EO-pretrained ResNet-50 [15] and ViT-Small [7] backbones [28] to maintain comparability in pretraining data. We include TerraMind (ViT-Small) [17] as a reference, as its TerraMind-Base [18] variant achieved leading performance in the evaluations in the original NeuCo-Bench paper. SSL4EO-pretrained backbones are accessed via TorchGeo [26], while TerraMind weights are accessed via TerraTorch [11]. Embeddings are extracted with TerraTorch for consistent processing across models. For the SSL4EO-pretrained backbones, we consider the self-supervised methods DINO [5], MoCo [14], DECUR [29], SoftCon [30], MAE [13], and FGMAE [32]. Not all backbone-method combinations are evaluated due to limited availability in pretrained weights. To ensure comparability, all models are evaluated using Sentinel‑2 L1C inputs only, although TerraMind supports additional modalities.
@@ 3.3 Temporal and Spatial Aggregation
The considered GeoFM encoders produce dense spatial representations: ResNet outputs feature maps of shape $C\times H\times W$, while ViT models produce patch tokens of shape $N_{\text{patch}}\times D$. Image inputs are resized to $224\times 224$ to match pretraining resolution. To obtain a 1D fixed‑size embedding per sample, we apply global mean pooling, across spatial dimensions for ResNet and across tokens for the ViT. We additionally evaluate max and min pooling for both architectures, and for ViTs also analyze CLS‑token embeddings. Embeddings are evaluated at their native dimensionality (2048 for ResNet‑50 and 384 for ViT‑Small).
@@ 4 Experiments and Results
We evaluate embedding design choices for GeoFMs across backbone choice, pretraining strategy, spatial aggregation, and representation depth. We begin with global per-task trends before analyzing each design dimension.
@@ 4.1 General Per-Task Trends
Semantic land-cover targets. Land-cover Agriculture and Land-cover Forest achieve the highest median $R^{2}$ values (both above 0.80) with low variability, indicating that these signals are reliably captured across embeddings. Crops shows moderate spread. While peak $R^{2}$ values are partially saturated, the Q-score reveals meaningful differences in robustness not visible from mean accuracy alone.
Continuous biophysical and atmospheric targets are more challenging. Biomass and Heat Island Mean show lower median $R^{2}$, and Clouds exhibits strong sensitivity to representation design with wide variation across methods, indicating greater dependence on representation choice. Heat Island Std is the most difficult target overall.
@@
Semantic land‑cover targets. For Crops, Land-cover Agriculture, and Land-cover Forest, ResNets remain competitive and in some cases slightly exceed ViT performance. This suggests that global semantic composition signals are less sensitive to backbone choice, whereas geophysical and atmospheric variables benefit from richer, long‑range spatial modeling. Across all configurations, TerraMind (ViT‑Small with multimodal pretraining on TerraMesh [2]) achieves the strongest and most stable performance.
These trends align with prior NeuCo‑Bench backbone comparisons performed under matched embedding dimensionality [27], where CNN embeddings were downsampled and ViT embeddings padded. Together with our results, this indicates that the observed ResNet-ViT gap cannot be attributed to embedding dimensionality or resizing artifacts.
@@ 4.3 Task-Specific Effects of SSL Objectives
These trends are consistent across both $R^{2}$ and Q-score. While $R^{2}$ reflects accuracy, the Q-score magnifies robustness differences, and in our results leads to clearer separation between methods. Objective-dependent differences are also more pronounced for transformer backbones. For ResNet models, variation across SSL objectives is weaker, as all methods struggle on continuous biophysical and atmospheric targets. Overall, pretraining objective shapes task-specific representation profiles. Contrastive objectives (e.g., DINO) favor semantic composition signals, whereas reconstruction-based objectives (e.g., MAE/FGMAE) better capture continuous geophysical variation. Rather than yielding a single universally best method, SSL objectives produce complementary strengths, an observation we will revisit in the concatenation experiments.
@@ 4.4 Impact of Spatial Pooling
Figure 3: Per-task Q-score comparison of spatial aggregation method for ResNet-50 (left) and ViT-Small (right). We use final-layer embeddings with mean, min, or max pooling (or the CLS token for ViT) and average scores across models; negative scores are clipped to zero. For ResNet, mean pooling performs best across tasks, with max pooling outperforming min pooling. For ViT, mean pooling again performs best, with CLS comparable on most tasks, while min and max pooling are similar but weaker, especially on continuous biophysical targets. A corresponding task $R^{2}$ plot is provided in the supplementary material.
@@ 4.5 Complementarity Through Concatenation
We analyze embedding concatenation from two perspectives: (i) per-task gains relative to the stronger task-specific baseline, and (ii) overall gains relative to the strongest baseline across tasks. These two measurements capture distinct effects. When concatenating embeddings under the multi‑task NeuCo‑Bench setting, two effects are possible. First, if individual baselines excel on different tasks, concatenation can improve the overall score simply by retaining the stronger signal per task, even when per‑task gains remain small or negative. Second, if the representations encode genuinely complementary information, concatenation may also yield per‑task improvements beyond both baselines, resulting in gains at both levels.
@@
Across DINO, MAE, and SoftCon, intra‑method concatenation yields only modest improvements (below $\sim$0.03 R2 and below 1 Q‑score point overall), with consistently small per‑task gains (Figure 4). This matches our earlier observation that Mean and CLS embeddings from the same ViT model show highly aligned task profiles and similar overall performance. Given this overlap, few complementary task scores can be exploited for overall improvements.
@@ 4.6 Strength of Intermediate CNN Layers
For ViT‑Small, performance increases over the first few layers and then saturates. Most SSL variants reach near‑peak accuracy by layers 3–5, with only minor changes afterward. Semantic land‑cover targets benefit slightly from deeper layers, whereas continuous biophysical and atmospheric targets (Biomass, Clouds, Heat Island) remain relatively stable across depth and, in some cases (e.g., DINO), degrade modestly in the final layers. This suggests that early transformer layers already capture the information needed for aggregated environmental prediction, while deeper layers increasingly emphasize semantic abstraction.
ResNet‑50, in contrast, exhibits a pronounced inverted‑U pattern. Both $R^{2}$ and Q‑score peak at intermediate stages (layers 2–4) and decline at the final stage. The drop is primarily driven by the continuous biophysical targets, which improve substantially at intermediate layers, reaching performance competitive with ViT representations, before degrading in the deepest layer. Semantic land‑cover targets continue to improve more steadily with depth. A resized final‑layer reference confirms that this effect is not attributable to embedding dimensionality.
Overall, optimal layer depth is task‑dependent. While final‑layer embeddings are commonly used by default, intermediate layers, especially for CNN backbones, often provide stronger and more stable representations for aggregated biophysical targets. These trends hold consistently across both $R^{2}$ and Q‑score.